# Community-Based Early-Stage Chronic Kidney Disease Screening using Explainable Machine Learning for Low-Resource Settings

### v1

@@ Abstract
Background: Early detection of chronic kidney disease (CKD) is essential for preventing progression to end-stage renal disease. However, existing screening tools—primarily developed using populations from high-income countries—often underperform in Bangladesh and South Asia, where risk profiles differ. Most of these tools rely on simple additive scoring functions and are based on data from patients with advanced-stage CKD. Consequently, they fail to capture complex interactions among risk factors and are limited in predicting early-stage CKD.
Objective: Develop and evaluate an explainable machine learning (ML) framework for community-based early-stage CKD screening for low-resource settings, tailored to the Bangladeshi and South Asian population context.
Methods: We used a community-based dataset from Bangladesh, the first such CKD dataset in South and South Asia, and evaluated twelve ML classifiers across multiple feature domains. Ten complementary feature selection techniques were applied to identify robust, generalizable predictors. The final models were assessed using 10-fold cross-validation. External validation was conducted on three independent datasets from India, the UAE, and Bangladesh. SHAP (SHapley Additive exPlanations) was used to provide model explainability.
Results: An ML model trained on an RFECV-selected feature subset achieved a balanced accuracy of 90.40%, whereas minimal non-pathology-test features demonstrated excellent predictive capability with a balanced accuracy of 89.23%, often outperforming larger or full feature sets. Compared with existing screening tools, the proposed models achieved substantially higher accuracy and sensitivity while requiring fewer and more accessible inputs. External validation confirmed strong generalizability with 78% to 98% sensitivity. SHAP interpretation identified clinically meaningful predictors consistent with established CKD risk factors.
Conclusions: Accurate, interpretable, and scalable early-stage CKD screening is achievable using only non-pathology-test features. This framework offers a practical pathway for community-level CKD screening in resource-constrained settings where laboratory testing is limited.
@@ 1 Introduction
The kidneys play a vital role in maintaining physiological balance within the human body. Their primary functions include filtering waste products and excess fluids from the blood, regulating blood pressure through renin-angiotensin pathways, maintaining acid-base balance, managing electrolyte levels, and stimulating red blood cell production through the secretion of erythropoietin [lote1994principles]. Chronic kidney disease (CKD) is a progressive condition marked by the gradual decline in kidney function, eventually requiring renal replacement therapy such as dialysis or kidney transplantation. Clinically, CKD is diagnosed by persistent markers of kidney damage or a sustained reduction in the estimated glomerular filtration rate (eGFR) below 60 mL/min/1.73 $m^{2}$ for more than three months [levin2013kidney].
@@
CKD represents a critical and growing global health challenge, with profound socioeconomic implications. Globally, kidney disease affects approximately 850 million people [jager2019single], with CKD estimated to have a prevalence of 11–13% [hill2016global]. In 2017, around 1.2 million people died from CKD and it was the 12th leading cause of death worldwide [bikbov2020global]. The overall number of CKD-related deaths has increased substantially over the recent decade, yet resources for managing this silent epidemic remain limited in most developing regions [bikbov2020global]. In low- and middle-income countries (LMICs), including South Asia, the burden is particularly severe due to limited healthcare infrastructure, shortage of nephrology care, and poor access to diagnostics [abraham2016chronic]. In Bangladesh, community-based studies in urban and rural populations suggest a prevalence rate of CKD 26% and 22%, respectively [anand2014high, sarker2021community, banik2021prevalence], with most patients unaware of their condition [iqbal2018knowledge].
@@
Early-stage CKD detection is essential to delay or prevent the onset of end-stage kidney disease (ESKD), reduce cardiovascular complications, and avoid the substantial costs associated with renal replacement therapies (RRT). Evidence suggests that CKD progression can be slowed considerably through early lifestyle modification, pharmacological management of comorbidities, and patient education [pollock2024framework]. However, early-stage CKD is typically asymptomatic; only about 4% of individuals with CKD stages 1–2 and fewer than 10% with stage 3 are aware of their condition [plantinga2008patient], and most patients are diagnosed only after clinical symptoms emerge, by which time renal impairment is often irreversible [levey2022chronic]. The failure to detect CKD at an early stage not only worsens patient outcomes but also imposes a significant economic burden on health systems [luyckx2021global], particularly in low-income settings where RRT is prohibitively expensive and largely inaccessible [niang2018hemodialysis].
@@
Standard clinical diagnosis of CKD relies on laboratory or pathology tests such as serum creatinine levels, estimated glomerular filtration rate (eGFR), and the urinary albumin-to-creatinine ratio (ACR) [levin2024executive]. These assessments require access to well-equipped diagnostic facilities, trained personnel, and regular follow-up—resources that are scarce in many low-resource settings [stanifer2014chronic]. Early diagnosis is particularly challenging in the Global South, including Bangladesh, where a large proportion of the population lives in rural and peri-urban communities with limited access to healthcare infrastructure, diagnostic laboratories, and nephrology specialists [kabir2023capacity]. In addition, poor awareness among patients and providers, inadequate communication between them, low risk perception among high-risk individuals, and the financial burden associated with travel, laboratory testing, and clinical consultations further deters early CKD diagnosis [zeba2020early, jafar2020access, nazar2014barriers, kahn2015chronic]. Together, these structural and socioeconomic barriers contribute to delayed recognition of CKD and missed opportunities for timely intervention.
@@
Several community-based CKD screening tools have been developed over the past two decades to facilitate risk assessment [stolpe2022external]; however, most were designed in high-income countries and exhibit limited relevance for low-resource settings (Table 2 provides a summary). Prominent tools include the SCORED [bang2007development] from the United States , kshirsagar2008simple based on the U.S. population, the thakkinstian2011simple developed in Thailand, the kearns2013predicting from England, and kwon2019simple from South Korea. These tools rely largely on demographic and medical history variables such as age, sex, diabetes, hypertension, and cardiovascular disease, with several incorporating clinical examination features (e.g., anemia, blood pressure, dipstick proteinuria). Importantly, most were designed to detect CKD stages 3–5, thereby overlooking early-stage disease where intervention is most impactful. Their dependency on Western cohorts also limits applicability to populations in the Global South, where environmental exposures, health system constraints, and sociodemographic risk patterns differ markedly. Moreover, most models employ simple additive scoring functions, restricting their ability to capture complex interactions among risk factors. These limitations underscore the need for context-appropriate, low-cost, and explainable machine learning approaches tailored to early-stage CKD screening in resource-constrained environments.
@@
Machine learning (ML)–based predictive tools, particularly those relying on easily obtainable features such as age, blood pressure, and medical history, offer a promising avenue for improving CKD risk assessment in low-resource settings. A substantial number of studies have applied ML techniques for CKD identification, and comprehensive summaries of these works are available in several review articles [sanmarchi2023predict, delrue2024application, gogoi2025machine]. Notably, the vast majority of ML studies are based on the UCI-2015 CKD dataset [chronic_kidney_disease_336], which we also used for external validation [gogoi2025machine]. However, many of these studies include pathology-test-based markers such as serum creatinine or eGFR as input features for model development (e.g., [gogoi2025chronic, dharmarathne2024diagnosis, islam2023chronic, ghosh2025machine, ghosh2024investigation]). This practice introduces a form of information leakage and circular reasoning, as these biomarkers are themselves the clinical criteria used to diagnose CKD; if such values are already available, an ML model becomes redundant. Moreover, pathology tests such as serum creatinine and eGFR are often inaccessible in rural or peri-urban areas and remain financially prohibitive for many individuals in the Global South. Consequently, these ML studies are unsuitable for community-based CKD screening and cannot be deployed in resource-limited populations. To the best of our knowledge, no prior ML study has focused specifically on community-based early-stage CKD screening while performing extensive feature selection, rigorous external validation across multiple datasets, and systematic comparison with established clinical screening tools as conducted in this work.
@@ 2 Methodology
Figure 1 provides an overview of our methodological pipeline. The workflow consists of several steps: (i) dataset preprocessing, including imputation, discretization, harmonization and one-hot encoding; (ii) feature ranking and selection using statistical, regularization-based, and wrapper-based methods; (iii) model development through training a diverse set of machine learning classifiers; (iv) performance evaluation, incorporating stratified cross-validation, benchmarking against state-of-the-art (SOTA) methods, and external validation on independent datasets; and (v) model explainability using SHAP to interpret both global and individual-level predictions. Together, these steps establish a rigorous and interpretable framework for early CKD prediction.
@@ 2.1 Datasets
The dataset used in this study originates from a community-based CKD screening conducted in the Mirzapur sub-district of Tangail, Bangladesh, a rural and peri-urban region covered by the Mirzapur demographic surveillance system (DSS). Adults aged $\geq 18$ years with at least five years of residency were selected using age-stratified random sampling, yielding 928 enrolled participants, of whom 872 completed all study procedures. Following home-based enrolment, trained community health workers obtained written informed consent, performed interviews and physical examinations, and referred participants to a hospital laboratory for the first assessment, during which blood and urine samples were collected to measure serum creatinine, fasting glucose, lipid profile, hemoglobin, and the urine albumin-to-creatinine ratio (ACR). Participants exhibiting an estimated glomerular filtration rate (eGFR) <60<60 mL/min/1.73 m2 and/or ACR $\geq 30$ mg/g were invited for a second assessment after a 3-month interval to confirm persistent abnormalities. CKD diagnosis was made using the CKD-EPI equation to estimate eGFR and stage following the National Kidney Foundation Kidney Disease Outcomes Quality Initiative (NKF/KDOQI) guidelines [inker2014kdoqi], requiring reduced eGFR and/or albuminuria sustained for at least three months. The study received ethics approval from the ICDDR,B Research Review Committee and Ethical Review Committee (registration no. 19081), and all participants provided written informed consent. Further details on the data collection process are provided in [sarker2021community].
@@
In this study, we focused exclusively on participants with CKD stages 1 to 3, as these stages represent the early phase of kidney impairment when symptoms are typically absent and community-based screening can still play a critical preventive role. Participants in stages 4 and 5 were excluded because these advanced stages are comparatively rare in the dataset and are usually accompanied by clinically evident symptoms, making them less relevant to early screening. After removing records with missing values, the final analytic dataset comprised 284 participants, including 112 CKD (stage 1: n = 25; stage 2: n = 51; stage 3: n = 36), and 172 non-CKD controls. Table 1 presents the descriptive characteristics of the 24 study variables, all of which were transformed into categorical variables for analysis. These variables span multiple groups, including socio-demographic (SD) factors (age group, gender, literacy status, occupation, and marital status), lifestyle and habit (LH) variables (sleep duration, tobacco smoking, and smokeless tobacco use), medical history (MH) (hypertension, diabetes, heart disease, stroke, and family histories of major conditions), and physical or clinical examination (CE) (body mass index (BMI) category, abdominal obesity, undernutrition status, and anemia), and pathology tests (Path) (presence of urinary red blood cells, low serum albumin, hypercholesterolemia, low HDL-cholesterol, and hypertriglyceridemia). These variables collectively represent the multidimensional risk profile relevant to early-stage CKD screening.
Table 1: Primary dataset description
@@
For external validation, we incorporated three publicly available datasets (summarized in Table 1). The UCI-2015 CKD dataset [chronic_kidney_disease_336], collected from a hospital in India, is the most widely used benchmark dataset for machine learning research on CKD detection. It contains clinical records from 400 patients admitted over a two-month period, comprising 250 CKD cases and 150 non-CKD controls. The dataset includes 24 features spanning multiple groups: one socio-demographic variable, one lifestyle and habit-related variable, four medical history variables, four clinical examination measures, and fourteen pathology test results, including two standard pathology tests for CKD diagnosis, such as urinary albumin and serum creatinine. Notably, CKD stages are not reported in this dataset.
Another dataset used for external validation is the UCI-2023 CKD dataset [risk_factor_prediction_of_chronic_kidney_disease_857], collected at Enam Medical College in Savar, Bangladesh. It contains records from 200 patients, including 128 confirmed CKD cases and 72 non-CKD controls, with CKD stages 1–5 explicitly annotated. The dataset comprises 27 features spanning several groups: one socio-demographic attribute, one lifestyle and habit-related variable, four medical history indicators, six clinical examination measurements, and fifteen pathology test results, including three standard pathology tests for CKD diagnosis, such as urinary albumin, serum creatinine, and eGFR.
The third dataset is the TH CKD dataset [Al-Shamsi2018], derived from hospital records at Tawam Hospital in Al-Ain City, Abu Dhabi, and is one of the few Middle Eastern clinical datasets used for CKD risk modeling. It includes 491 adult patients, of whom 56 were identified with moderate-to-advanced CKD (stages 3–5). The remaining 435 participants were labeled as the “non-CKD” category. However, because early-stage CKD (stages 1–2) was not distinguished from healthy controls, this negative class does not constitute a clinically reliable non-CKD population. Consequently, these samples were excluded from cross-validation in our study, and only the confirmed CKD cases were used for external evaluation. The dataset contains 21 features spanning multiple groups: two socio-demographic variables, one lifestyle and habit-related factor, ten medical history attributes, three clinical examination measures, and five pathology test indicators, including two standard pathology tests for CKD diagnosis, such as serum creatinine and eGFR.

### v2

@@ Abstract
Background: Early detection of chronic kidney disease (CKD) is essential for preventing progression to end-stage renal disease. However, existing screening tools—primarily developed using populations from high-income countries—often underperform in Bangladesh and South Asia, where risk profiles differ. Moreover, many machine learning–based CKD studies rely heavily on pathology-test variables and are developed using hospital-based datasets, limiting their applicability for community-level screening and for settings where laboratory testing is not readily accessible.
Objective: To develop and evaluate an explainable machine learning (ML) framework for community-based early-stage CKD screening that derives and evaluates optimized predictor subsets from both all available variables and variables excluding pathology tests, enabling accurate and practical risk assessment in low-resource settings.
Methods: A community-based CKD dataset from Bangladesh was used to develop predictive models. Variables were organized into clinically meaningful feature groups, and ten complementary feature selection methods were applied to identify robust predictor subsets. Twelve ML classifiers were evaluated using nested cross-validation. Model performance was benchmarked against established CKD screening tools and externally validated on three independent datasets from India, the UAE, and Bangladesh. SHAP (SHapley Additive exPlanations) was used to interpret model predictions.
Results: The best-performing model using the optimized feature subset derived from all variables achieved a balanced accuracy of 90.4%, while the subset derived from variables excluding pathology tests achieved comparable performance (89.23%), demonstrating the feasibility of accurate CKD risk prediction using minimal non-laboratory features. The proposed approach outperformed existing screening tools while requiring fewer and more accessible inputs. External validation demonstrated strong generalizability, with sensitivities ranging from 78% to 98%. SHAP analysis identified clinically meaningful predictors consistent with established CKD risk factors.
Conclusions: Accurate, interpretable, and scalable early-stage CKD screening is achievable using only non-pathology-test features. This framework demonstrates the potential for community-level CKD screening in resource-constrained settings.
@@ 1 Introduction
Chronic kidney disease (CKD) represents a critical and growing global health challenge, with profound socioeconomic implications. Globally, kidney disease affects approximately 850 million people [37], with CKD estimated to have a prevalence of 11–13% [26]. In 2017, around 1.2 million people died from CKD, and it was the 12th leading cause of death worldwide [9]. The overall number of CKD-related deaths has increased substantially over the recent decade, yet resources for managing this silent epidemic remain limited in most developing regions [9]. In low- and middle-income countries (LMICs), including South Asia, the burden is particularly severe due to limited healthcare infrastructure, shortage of nephrology care, and poor access to diagnostics [1]. In Bangladesh, community-based studies in urban and rural populations suggest a prevalence rate of CKD 26% and 22%, respectively [5, 68, 7], with most patients unaware of their condition [33].
@@
Early-stage CKD detection is essential to delay or prevent the onset of end-stage kidney disease (ESKD), reduce cardiovascular complications, and avoid the substantial costs associated with renal replacement therapies (RRT). Evidence suggests that CKD progression can be slowed considerably through early lifestyle modification, pharmacological management of comorbidities, and patient education [61]. However, early-stage CKD is typically asymptomatic; only about 4% of individuals with CKD stages 1–2 and fewer than 10% with stage 3 are aware of their condition [60], and most patients are diagnosed only after clinical symptoms emerge, by which time renal impairment is often irreversible [48]. The failure to detect CKD at an early stage not only worsens patient outcomes but also imposes a significant economic burden on health systems [51], particularly in low-income settings where RRT is prohibitively expensive and largely inaccessible [57].
@@
Standard clinical diagnosis of CKD relies on laboratory or pathology tests such as serum creatinine levels, estimated glomerular filtration rate (eGFR), and the urinary albumin-to-creatinine ratio (ACR) [49]. These assessments require access to well-equipped diagnostic facilities, trained personnel, and regular follow-up—resources that are scarce in many low-resource settings [69]. Early diagnosis is particularly challenging in the Global South, including Bangladesh, where a large proportion of the population lives in rural and peri-urban communities with limited access to healthcare infrastructure, diagnostic laboratories, and nephrology specialists [39]. In addition, poor awareness among patients and providers, inadequate communication between them, low risk perception among high-risk individuals, and the financial burden associated with travel, laboratory testing, and clinical consultations further deters early CKD diagnosis [74, 36, 56, 40]. Together, these structural and socioeconomic barriers contribute to delayed recognition of CKD and missed opportunities for timely intervention.
@@
Several community-based CKD screening tools (e.g., SCORED [6]) have been developed over the past two decades to facilitate risk assessment [70]; however, most were designed in high-income countries and exhibit limited relevance for low-resource settings (Table 1 provides a summary). Prominent tools include the SCORED [6] from the United States, Kshirsagar et al. [46] based on the U.S. population, the Thakkinstian et al. [71] developed in Thailand, the Kearns et al. [42] from England, and Kwon et al. [47] from South Korea. These tools rely largely on demographic and medical history variables such as age, sex, diabetes, hypertension, and cardiovascular disease, with several incorporating clinical examination features (e.g., anemia, blood pressure, dipstick proteinuria). Importantly, most were designed to detect CKD stages 3–5, thereby overlooking early-stage disease where intervention is most impactful. Their dependency on Western cohorts also limits applicability to populations in the Global South, where environmental exposures, health system constraints, and sociodemographic risk patterns differ markedly. Moreover, most models employ simple additive scoring functions, limiting their ability to capture complex interactions among risk factors. These limitations underscore the need for context-appropriate, low-cost, and explainable machine learning approaches tailored to early-stage CKD screening in resource-constrained environments.
@@
Machine learning (ML)–based predictive tools, particularly those relying on easily obtainable features such as age, blood pressure, and medical history, offer a promising avenue for improving CKD risk assessment in low-resource settings. A substantial number of studies have applied ML techniques for CKD identification, and comprehensive summaries of these works are available in several review articles [67, 16, 24, 66, 44]. Notably, the vast majority of ML studies are based on the UCI-2015 CKD dataset [64], which we also used for external validation. However, many of these studies include pathology-test-based markers such as serum creatinine or eGFR as input features for model development (e.g., [17, 34, 62, 23, 8, 38, 58]). This practice introduces a form of information leakage and circular reasoning, as these biomarkers are themselves the clinical criteria used to diagnose CKD; if such values are already available, an ML model becomes redundant. Moreover, pathology tests such as serum creatinine (SC) and eGFR are often inaccessible in rural or peri-urban areas and remain financially prohibitive for many individuals in the Global South. Consequently, these ML studies may be less suitable for community-based CKD screening and may have limited applicability in resource-limited populations where access to laboratory testing is constrained. While previous machine learning studies have investigated CKD prediction, none have focused specifically on community-based early-stage CKD screening while integrating extensive feature selection, systematic comparison with existing screening tools, and external validation across multiple datasets.
Building on these observations, this study proposes a machine learning–based framework for community-level early-stage CKD screening that emphasizes clinically accessible features, systematic feature selection, and rigorous validation across multiple datasets. This study makes several distinct contributions to the CKD screening literature:
  • 1.
    It develops a community-based early-stage CKD screening framework using population-level data from Bangladesh, focusing specifically on CKD stages 1–3, where early intervention is most beneficial.
  • 2.
    It systematically investigates the predictive utility of clinically meaningful feature groups (socio-demographic, lifestyle and habit, medical history, clinical examination, and pathology tests) and their combinations, reflecting their impact on CKD screening.
  • 3.
    It employs a rigorous feature selection framework using ten complementary methods to identify stable and clinically meaningful predictors.
  • 4.
    It derives and evaluates optimized feature subsets from (i) all available variables and (ii) variables excluding pathology tests, thereby identifying minimal yet effective predictor sets suitable for community-level screening in low-resource settings.
  • 5.
    It performs systematic benchmarking against established CKD screening tools on the primary dataset using machine learning models based on the derived feature subsets, demonstrating improved predictive performance.
  • 6.
    It conducts cross-dataset external validation by training models on the primary dataset using the derived feature subsets and evaluating them on multiple independent datasets, demonstrating the robustness and generalizability of these feature subsets across diverse populations.
2 Related work
Recent advances in machine learning (ML) have led to a growing body of research on predicting chronic kidney disease (CKD) using structured health data. Several recent surveys summarize the application of artificial intelligence techniques in CKD-related tasks, including disease detection, risk prediction, prognosis, and monitoring [67, 16, 24, 66, 44]. Collectively, these reviews highlight the potential of ML to support CKD research and clinical decision-making. At the same time, they identify recurring limitations in the existing literature, including limited dataset diversity, insufficient focus on early-stage CKD detection, and challenges in translating high-performing models into clinically meaningful and practically deployable screening tools. Table 1 summarizes representative recent ML-based CKD prediction studies most relevant to the scope of this work. The comparison highlights differences across studies in terms of research objectives (e.g., community-level screening, early-stage detection, or applications in low-resource settings), the types of features considered (including socio-demographic, lifestyle, clinical examination, and pathology variables), and the extent to which models rely on pathology-test or gold-standard clinical biomarkers such as serum creatinine or eGFR. It also contrasts key methodological aspects, including feature selection strategies, explainability techniques, and validation protocols. This structured comparison helps clarify how the present work differs from prior studies and motivates the methodological choices adopted in this study.
Table 1: Summary of recent ML-based studies on CKD prediction
Objective Considered features Approach
Study Year Dataset
Community-level
Early-stage screening
Low-resource settings
Socio-demographic
Lifestyle and habit
Medical history
Clinical exam
Pathology test
Total
Used SC or eGFR
FS among all features
FS excluding pathology
Explainability
External validation
[15] 2022 SPH $\sim$ 2 0 3 3 10 18
[75] 2023 TH $\sim$ 2 2 5 3 3 16
[22] 2024 TH $\sim$ 2 2 9 3 5 22
[43] 2025 Bab Al-Rayan 1 0 2 0 7 10
[31] 2023 Buner Medical 2 0 0 1 17 20
[30] 2025 Buner Medical 2 0 0 1 12 15
[29] 2026 Buner Medical 2 0 0 1 17 20
[17] 2024 UCI-2015 1 0 2 3 4 10
[34] 2023 UCI-2015 0 0 0 3 5 8
[62] 2023 UCI-2015 1 0 4 3 10 18
[23] 2025 UCI-2015 1 1 4 4 13 23
[8] 2025 UCI-2015 1 1 4 4 14 24
[38] 2025 UCI-2015 1 0 2 1 2 6
[58] 2025 UCI-2015 1 1 4 3 13 22
[54] 2025 UCI-2015 1 1 4 4 14 24
[55] 2025 UCI-2015 1 0 2 2 7 11
[28] 2025 UCI-2015, UCI-2023 $\sim$ 1 0 4 4 14 23
[63] 2024 UCI-2015, UCI-2023 $\sim$ 1 1 3 3 12 20
[53] 2023 UCI-2015, UCI-2023 $\sim$ 1 1 4 1 6 13
This study TangailBD, 5 3 7 4 5 24
(TH, UCI-2015, UCI-2023)
FS: Feature Selection, SC: Serum Creatinine, –: unknown as CKD stages not specified,
$\sim$: partially, as the dataset includes both early and advanced CKD stages
Most recent ML-based CKD prediction studies [17, 34, 62, 23, 8, 38, 58, 54, 55, 75, 22] rely heavily on three publicly available datasets, most notably the earlier UCI-2015 dataset [64] and the more recent UCI-2023 [35] and TH [4] datasets. A few studies have developed CKD prediction models using their own datasets, such as SPH [15], Bab Al-Rayan [43], and Buner Medical [31] datasets. However, these datasets are primarily collected in hospital settings and provide limited information on socio-demographic characteristics, lifestyle factors, and behavioral habits, which are important determinants for community-level risk assessment. In addition, these datasets offer limited support for early-stage CKD screening. For example, the UCI-2015 [64] dataset does not provide CKD stage information for individual subjects, while the TH dataset [4] primarily covers patients with CKD stages 3–5 without identifying stages at the individual level. Only the UCI-2023 [35] dataset includes subject-level CKD stage annotations, which are necessary for stage-aware prediction tasks. Similarly, the SPH dataset used by Debal and Sitote [15] reports that it includes CKD stages 1–5, but detailed stage-level annotations are not clearly described. The studies used Buner Medical dataset [31, 30, 29] and Bab Al-Rayan dataset [43] also do not explicitly specify CKD stages. As a result, these datasets have limited applicability for developing models specifically aimed at community-based early-stage CKD screening.
Furthermore, the majority of existing studies [17, 34, 62, 23, 8, 38, 58, 15, 75, 22, 43] incorporate pathology-test variables such as serum creatinine (SC) or estimated glomerular filtration rate (eGFR) as input features during model development. Because these biomarkers are themselves part of the clinical diagnostic criteria for CKD, models that rely on them may provide limited additional value for early risk screening. In addition, such pathology tests are often unavailable or costly in rural or peri-urban settings and may require laboratory infrastructure that is not readily accessible in many parts of the Global South. Consequently, models built around these variables may be less suitable for community-based CKD screening and may have limited practical applicability in resource-constrained environments where laboratory testing is not routinely available.
Few studies [31, 30, 29, 54, 55] avoid using SC or eGFR as predictors; however, their primary focus is not on developing screening models for low-resource settings. Consequently, they do not explicitly investigate feature subsets excluding pathology-test variables or evaluate models designed to operate without such inputs. In addition, none of these studies reports external validation, which limits the ability to assess the generalizability of the proposed models across different populations.
From a methodological perspective, several studies [34, 62, 23, 8, 58, 15, 75, 43, 30, 55, 28, 63, 53] have explored feature selection techniques to identify important predictors for CKD classification. However, most of these studies focus on selecting optimal features from the full set of available variables, which often include pathology-test biomarkers such as SC or eGFR. While feature selection can improve model performance and reduce dimensionality, existing studies generally do not explicitly investigate feature subsets that exclude pathology-test variables. This distinction is particularly important for community-level screening in low-resource environments, where such laboratory measurements may not be readily available. As a result, prior work provides limited insight into whether reliable CKD prediction can be achieved using only clinically accessible features.
Several studies (e.g., [17, 23, 8, 38, 58, 75, 22, 30, 28]) have also increasingly incorporated explainable artificial intelligence (XAI) techniques to improve model interpretability. Methods such as SHAP or feature importance analysis have been used to identify influential predictors and provide insights into model behavior. While these approaches contribute to transparency and clinical interpretability, they apply explainability primarily as a post hoc analysis without linking it to the broader objective of developing screening models for community-level deployment.
Another important methodological aspect concerns model validation. Only a few studies [28, 63, 53] attempt cross-dataset validation by training models on the UCI-2015 dataset and testing them on the UCI-2023 dataset. However, this approach still inherits the limitations associated with both datasets, including their hospital-based nature and dependence on pathology-test variables. In addition, some of these studies [28, 63] still incorporate SC or eGFR as input features, which reduces their applicability for screening in resource-constrained environments. Consequently, rigorous external validation of models designed specifically for community-level CKD screening—particularly those relying on non-pathology features—remains largely unexplored.
Overall, the existing ML-based CKD prediction literature shows several limitations for community-level early-stage screening. Most studies rely on hospital-based datasets that lack comprehensive socio-demographic and lifestyle information relevant for population-level risk assessment. Many models also depend on pathology-test variables such as serum creatinine or eGFR, which are themselves diagnostic markers and may therefore provide limited additional value for predictive modeling; moreover, these tests may be unavailable in low-resource settings. While some studies apply feature selection and explainability techniques, they rarely investigate predictor subsets that exclude pathology-test variables. In addition, rigorous external validation across independent populations remains limited. These gaps highlight the need for ML frameworks that rely on clinically accessible features and are specifically designed and validated for community-based early-stage CKD screening.
@@ 3 Methodology
Figure 1 provides an overview of the methodological pipeline used in this study. The workflow consists of several stages: (i) dataset preprocessing, including imputation, discretization, feature harmonization, and one-hot encoding to ensure consistency across the primary dataset and the external datasets; (ii) feature ranking and selection performed exclusively on the primary dataset (TangailBD) using statistical, regularization-based, and wrapper-based techniques to identify informative predictor subsets suitable for community-level early-stage CKD screening in low-resource settings; (iii) using these derived feature subsets, multiple machine learning classifiers are trained and evaluated on the primary dataset together with hyperparameter optimization conducted within a stratified nested cross-validation framework, where model performance was estimated using an outer stratified 10-fold cross-validation while hyperparameters were tuned within each training fold using an inner 5-fold cross-validation; (iv) the best-performing models are further assessed through benchmarking against established CKD screening tools on the primary dataset using the same derived feature subsets; (v) external validation by applying the best-performing models trained on the primary dataset to three independent datasets (UCI-2015, UCI-2023, and TH) to assess robustness and generalizability of the derived feature subsets across different populations; and (vi) model explainability using SHAP (SHapley Additive exPlanations) to interpret both global feature importance and individual-level predictions.
@@ 3.1 Datasets
The dataset used in this study originates from a community-based CKD screening conducted in the Mirzapur sub-district of Tangail, Bangladesh, a rural and peri-urban region covered by the Mirzapur demographic surveillance system (DSS). Adults aged $\geq 18$ years with at least five years of residency were selected using age-stratified random sampling, yielding 928 enrolled participants, of whom 872 completed all study procedures. Following home-based enrolment, trained community health workers obtained written informed consent, performed interviews and physical examinations, and referred participants to a hospital laboratory for the first assessment, during which blood and urine samples were collected to measure serum creatinine, fasting glucose, lipid profile, hemoglobin, and the urine albumin-to-creatinine ratio (ACR). Participants exhibiting an estimated glomerular filtration rate (eGFR) <60<60 mL/min/1.73 m2 and/or ACR $\geq 30$ mg/g were invited for a second assessment after a 3-month interval to confirm persistent abnormalities. CKD diagnosis was made using the CKD-EPI equation to estimate eGFR and stage following the National Kidney Foundation Kidney Disease Outcomes Quality Initiative (NKF/KDOQI) guidelines [32], requiring reduced eGFR and/or albuminuria sustained for at least three months. The study received ethics approval from the ICDDR,B Research Review Committee and Ethical Review Committee (registration no. 19081), and all participants provided written informed consent. Further details on the data collection process are provided in [68].
@@
In this study, we focused exclusively on participants with CKD stages 1 to 3, as these stages represent the early phase of kidney impairment when symptoms are typically absent and community-based screening can still play a critical preventive role. Participants in stages 4 (n=5) and 5 (n=1) were excluded because these advanced stages are comparatively rare in the dataset and are usually accompanied by clinically evident symptoms, making them less relevant to early screening. After removing records with missing values, the final analytic dataset comprised 284 participants, including 112 CKD (stage 1: n = 25; stage 2: n = 51; stage 3: n = 36), and 172 non-CKD controls. Table 2 presents the descriptive characteristics of the 24 study variables, all of which were transformed into categorical variables for analysis. These variables span multiple groups, including socio-demographic (SD) factors (age group, gender, literacy status, occupation, and marital status), lifestyle and habit (LH) variables (sleep duration, tobacco smoking, and smokeless tobacco use), medical history (MH) (hypertension, diabetes, heart disease, stroke, and family histories of major conditions), and physical or clinical examination (CE) (body mass index (BMI) category, abdominal obesity, undernutrition status, and anemia), and pathology tests (Path) (presence of urinary red blood cells, low serum albumin, hypercholesterolemia, low HDL-cholesterol, and hypertriglyceridemia). These variables collectively represent the multidimensional risk profile relevant to early-stage CKD screening.
Table 2: Primary dataset (TangailBD) description
@@
For external validation, we incorporated three publicly available datasets (summarized in Table 3).
Table 3: Summary of CKD datasets used in this study
Dataset Country Sample size CKD Stage Features
CKD Non-CKD Total SD LH MH CE Path Path* Total
TangailBD Bangladesh 112 172 284 1–3 5 3 7 4 5 5 29
UCI-2015 [64] India 250 150 400 1 1 3 5 12 2 24
UCI-2023 [35] Bangladesh 128 72 200 1–5 1 1 4 6 12 3 27
TH [4] UAE 56 56 3–5 2 1 10 3 3 2 21
*Gold-standard pathology tests for CKD diagnosis: serum creatinine, urinary albumin, urinary creatinine, eGFR, and uACR
The UCI-2015 CKD dataset [64], collected from a hospital in India, is the most widely used benchmark dataset for machine learning research on CKD detection. It contains clinical records from 400 patients admitted over a two-month period, comprising 250 CKD cases and 150 non-CKD controls. The dataset includes 24 features spanning multiple groups: one socio-demographic variable, one lifestyle and habit-related variable, four medical history variables, four clinical examination measures, and fourteen pathology test results, including two standard pathology tests for CKD diagnosis, such as urinary albumin and serum creatinine. Notably, CKD stages are not reported in this dataset.
Another dataset used for external validation is the UCI-2023 CKD dataset [35], collected at Enam Medical College in Savar, Bangladesh. It contains records from 200 patients, including 128 confirmed CKD cases and 72 non-CKD controls, with CKD stages 1–5 explicitly annotated. The dataset comprises 27 features spanning several groups: one socio-demographic attribute, one lifestyle and habit-related variable, four medical history indicators, six clinical examination measurements, and fifteen pathology test results, including three standard pathology tests for CKD diagnosis, such as urinary albumin, serum creatinine, and eGFR.
The third dataset is the TH CKD dataset [4], derived from hospital records at Tawam Hospital in Al-Ain City, Abu Dhabi, and is one of the few Middle Eastern clinical datasets used for CKD risk modeling. It includes 491 adult patients, of whom 56 were identified with moderate-to-advanced CKD (stages 3–5). The remaining 435 participants were labeled as the “non-CKD” category. However, because early-stage CKD (stages 1–2) was not distinguished from healthy controls, this negative class does not constitute a clinically reliable non-CKD population. Consequently, these samples were excluded from cross-validation in our study, and only the confirmed CKD cases were used for external evaluation. The dataset contains 21 features spanning multiple groups: two socio-demographic variables, one lifestyle and habit-related factor, ten medical history attributes, three clinical examination measures, and five pathology test indicators, including two standard pathology tests for CKD diagnosis, such as serum creatinine and eGFR.